Using MATLAB

3. Write a MATLA code to solve the following system of linear equations: 20y 3x 1z 3
Answer #1

Let's write the equations in simple form:

-x + y + 4z = -2

3x + 20y -z = 3

0x + 2y + 9z = 3

Following is the required code to solve the above problem.

% Matrix A contains the coefficients of the variables

A = [
-1 1 4
3 20 -1
0 2 9
]
  

% Matrix B contains the values at the Right Hand Side in the three equations
B = [ -2 ; 3 ; 3 ]

% compute the value for x, y, and z

solution = inv(A) * B

Here is a screenshot of the output for above code:

Default Term warning: function-/demo.m shadows a core library function -1 1 4 32?-1 2 9 2 solution = 3.29730 -0.32432 0.48541 sh-4.3$

Note: I was confused if the equations were correctly written. This is how they were in the question.

If you feel that the equations that i have written after simplifying doesn't match yours, please do COMMENT.

Hope this code meets all your requirements...

BEST WISHES!